2. Famous Cinema Theatres sell the movie tickets for $16 per ticket for evening screenings and sell 250 tickets a night on average. They estimate that the price elasticity of demand for tickets is (negative) 0.8.

Calculate the expected number of tickets sold if they reduce the ticket price to $14
1
Expert's answer
2019-05-27T11:58:33-0400
"Kd=\Delta Q/Q1:\Delta P/P1"

So,


"-0.8=Q2-250\/250:(14-16)\/16"

In this case the expected quantity is


"Q2=-0.8*(-1\/8)*250+250=275"

Q2=275 tickets